Numerical Analysis I, Fall 2020 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 11

1. Section 4.3: Problems 14, 19, 20, 22, 23, 24.

2. Section 4.7: Problems 2(b), 4(b), 11, 13, 14.